OpenStudy (emilyjones284):

Using u-sub Y/(y+1)^3 dy

OpenStudy (fibonaccichick666):

I'm assuming you have this \[\int\frac{1}{(y+1)^3}\]?

OpenStudy (emilyjones284):

No

OpenStudy (anonymous):

\[\int\frac{y}{(y+1)^3}~dy~~?\] Let \(u=y+1\), then \(y=u-1\) and \(du=dy\). \[\int\frac{y}{(y+1)^3}~dy=\int\frac{u-1}{u^3}~du\]